OpenStudy (anonymous):

Jeff is holding a deck of cards, and asks you to pick 2. What is the probability that you pick a heart the first time and a spade the second time? He says to put the first one back in the deck before picking the second one.

OpenStudy (anonymous):

\[\frac{\text{Number of Hearts}}{52} \times \frac{\text{Number of Spades}}{52} \implies \frac{13}{52} \times \frac{13}{52}\implies \space ?\]

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

Just multiply and you'll get you're answer.

OpenStudy (anonymous):

i gt 676

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

nvm i gt 0.25

OpenStudy (anonymous):

Look. \[\frac{13}{52} = \frac{1}{4}\] \[\frac 14 \times \frac 14 \implies \frac{1}{16}\]